# Gegenbauer Graph Neural Networks for Time-varying Signal Reconstruction

### v1

@@ III-C Reconstruction of Smooth Time-varying Graph Signals
where $\mathbf{J}\in{0,1}^{N\times M}$ denotes the sampling matrix, $\mathbf{Y}\in\mathbb{R}^{N\times M}$ represents the matrix of observed values, and $f(\mathbf{\tilde{X}})$ is a regularization function. From a machine learning standpoint, the first term in (4) corresponds to the MSE loss between the observed values and the reconstructed values, while the second term serves as a regularization component specifically tailored for time-varying signals.
@@ IV-B Gegenbauer Polynomials
The Gegenbauer polynomials $C_{k}^{(\alpha)}(z)$ of degree $k$ are orthogonal on the interval $z\in[-1,1]$ with respect to the weight function $(1-z^{2})^{\alpha-1/2}$ and are solutions to the Gegenbauer differential equation $(1-z^{2})y^{\prime\prime}-2(\mu+1)y^{\prime}+(v-\mu)(v+\mu+1)y=0$ [63]. Gegenbauer polynomials are a generalization of Chebyshev and Legendre polynomials to a $(2\alpha+2)$–dimensional vector space and are proportional to the ultraspherical polynomials $P_{k}^{(\lambda)}(z)$. We can represent Gegenbauer polynomials in terms of the Jacobi polynomials $J_{k}^{(\alpha,\beta)}(z)$ when $\alpha=\beta=\lambda-\frac{1}{2}$ by:
@@ V-B Datasets
Sea surface temperature dataset. The sea surface temperature dataset utilized in our study was obtained from the Earth System Research Laboratory  [^4] https://psl.noaa.gov/. This dataset consists of monthly measurements spanning from 1870 to 2014, with a spatial resolution of $1^{\circ}$ latitude ×$\times 1^{\circ}$ longitude. We specifically focused on the Pacific Ocean region, ranging from $170^{\circ}$ west to $90^{\circ}$ west and $60^{\circ}$ south to $10^{\circ}$ north. In accordance with [30], we randomly selected $100$ points within this region for analysis. The dataset covers a time period of $600$ months. The temperature data within the selected points range from $-0.01$ C to $30.72$ C, with an average value of $19.15$ C.
@@ V-D Experiments
For all datasets except for the SW06 experiment, we constructed graphs $G$ using the $k$-NN approach with the Euclidean distance between node locations and a Gaussian kernel as in [23]. Regarding the PM 2.5 concentration, the sampling densities $m$ are ${0.1,0.15,0.2,\ldots,0.45}$. For the sea-surface temperature, the sampling densities are specified as ${0.1,0.2,\ldots,0.9}$, while for the Intel Lab dataset, they are set to ${0.1,0.3,0.5,0.7}$. In all three datasets, we establish a connection between the graph’s nodes by setting $k=5$ in the $k$-NN algorithm. For the SW06 experiment, we also consider the sampling densities in the set ${0.1,0.2,\ldots,0.9}$. However, the SW06 experiment involved sensors placed in a three-dimensional (3D) environment, where the horizontal distances between sensors were in the order of kilometers and the vertical distances were in meters. In such a scenario, a $k$-NN approach alone does not capture the temperature fluctuations in the underwater sensor network. To address this issue, we applied a graph inference approach that tackles the Graph System Identification (GSI) problem using a regularized maximum likelihood (ML) criterion, as explained in Section III-D. This approach allows us to incorporate temperature fluctuations.
@@ V-F2 Time-dependency study
In this section, we conducted a thorough investigation to assess the effectiveness of the specialized loss function in capturing the temporal dependency of the dataset. To evaluate its impact, we performed experiments on two representative and challenging datasets: SW06 and Sea-Surface. These datasets were selected due to their real-world nature and the complex dynamics observed in open environments. In our analysis, we explored four different scenarios to compare the performance of our architecture. 1) We utilized the Sobolev smoothness term in its original form with $\lambda>0$. 2) We set the parameter $\epsilon$ in the Sobolev term to zero, effectively removing its influence. 3) We excluded the differential temporal operator $\mathbf{D}_{h}$ from the analysis. 4) We completely omitted the Sobolev smoothness regularization term and relied solely on the MSE as the error metric. The results of our experiments are summarized in Table IV. From the results obtained on the SW06 and Sea-surface datasets, we observe that the model utilizing the MSE and complete Sobolev term in its loss function outperforms the other variations. These findings imply that our regularization term for capturing time-dependency in the data is significantly enhancing the performance of the GegenGNN model. The specialized loss function addresses time-dependency by integrating the Sobolev smoothness term with the temporal difference operator $\mathbf{D}_{h}$. Notice that when using the Sobolev norm it is important to keep the $\epsilon$ term in Eq. (18) small to avoid perturbing the Laplacian matrix $\mathbf{L}$. The appropriate value for this term should be chosen via cross-validation methods.

### v2

@@ III-C Reconstruction of Smooth Time-varying Graph Signals
where $\mathbf{J}\in\{0,1\}^{N\times M}$ denotes the sampling matrix, $\mathbf{Y}\in\mathbb{R}^{N\times M}$ represents the matrix of observed values, and $f(\mathbf{\tilde{X}})$ is a regularization function. From a machine learning standpoint, the first term in (4) corresponds to the MSE loss between the observed values and the reconstructed values, while the second term serves as a regularization component specifically tailored for time-varying signals.
@@ IV-B Gegenbauer Polynomials
The Gegenbauer polynomials $C_{k}^{(\alpha)}(z)$ of degree $k$ are orthogonal on the interval $z\in[-1,1]$ with respect to the weight function $(1-z^{2})^{\alpha-1/2}$ and are solutions to the Gegenbauer differential equation $(1-z^{2})y^{\prime\prime}-2(\mu+1)y^{\prime}+(v-\mu)(v+\mu+1)y=0$ [63]. Gegenbauer polynomials are a generalization of Chebyshev and Legendre polynomials to a $(2\alpha+2)$–dimensional vector space and are proportional to the ultraspherical polynomials $P_{k}^{(\alpha)}(z)$. We can represent Gegenbauer polynomials in terms of the Jacobi polynomials $J_{k}^{(\lambda,\beta)}(z)$ when $\lambda=\beta=\alpha-\frac{1}{2}$ by:
@@ V-B Datasets
Sea-surface temperature dataset. The sea-surface temperature dataset utilized in our study was obtained from the Earth System Research Laboratory  [^4] https://psl.noaa.gov/. This dataset consists of monthly measurements spanning from 1870 to 2014, with a spatial resolution of $1^{\circ}$ latitude ×$\times 1^{\circ}$ longitude. We specifically focused on the Pacific Ocean region, ranging from $170^{\circ}$ west to $90^{\circ}$ west and $60^{\circ}$ south to $10^{\circ}$ north. In accordance with [30], we randomly selected $100$ points within this region for analysis. The dataset covers a time period of $600$ months. The temperature data within the selected points range from $-0.01$ C to $30.72$ C, with an average value of $19.15$ C.
@@ V-D Experiments
For all datasets except for the SW06 experiment, we constructed graphs $G$ using the $k$-NN approach with the Euclidean distance between node locations and a Gaussian kernel as in [23]. Regarding the PM 2.5 concentration, the sampling densities $m$ are $\{0.1,0.15,0.2,\ldots,0.45\}$. For the sea-surface temperature, the sampling densities are specified as $\{0.1,0.2,\ldots,0.9\}$, while for the Intel Lab dataset, they are set to $\{0.1,0.3,0.5,0.7\}$. In all three datasets, we establish a connection between the graph’s nodes by setting $k=5$ in the $k$-NN algorithm. For the SW06 experiment, we also consider the sampling densities in the set $\{0.1,0.2,\ldots,0.9\}$. However, the SW06 experiment involved sensors placed in a three-dimensional (3D) environment, where the horizontal distances between sensors were in the order of kilometers and the vertical distances were in meters. In such a scenario, a $k$-NN approach alone does not capture the temperature fluctuations in the underwater sensor network. To address this issue, we applied a graph inference approach that tackles the Graph System Identification (GSI) problem using a regularized maximum likelihood (ML) criterion, as explained in Section III-D. This approach allows us to incorporate temperature fluctuations.
@@ V-F2 Time-dependency study
In this section, we conducted a thorough investigation to assess the effectiveness of the specialized loss function in capturing the temporal dependency of the dataset. To evaluate its impact, we performed experiments on two representative and challenging datasets: SW06 and Sea-Surface temperature. These datasets were selected due to their real-world nature and the complex dynamics observed in open environments. In our analysis, we explored four different scenarios to compare the performance of our architecture. 1) We utilized the Sobolev smoothness term in its original form with $\lambda>0$. 2) We set the parameter $\epsilon$ in the Sobolev term to zero, effectively removing its influence. 3) We excluded the differential temporal operator $\mathbf{D}_{h}$ from the analysis. 4) We completely omitted the Sobolev smoothness regularization term and relied solely on the MSE as the error metric. The results of our experiments are summarized in Table IV. From the results obtained on the SW06 and Sea-surface temperature datasets, we observe that the model utilizing the MSE and complete Sobolev term in its loss function outperforms the other variations. These findings imply that our regularization term for capturing time-dependency in the data is significantly enhancing the performance of the GegenGNN model. The specialized loss function addresses time-dependency by integrating the Sobolev smoothness term with the temporal difference operator $\mathbf{D}_{h}$. Notice that when using the Sobolev norm it is important to keep the $\epsilon$ term in Eq. (18) small to avoid perturbing the Laplacian matrix $\mathbf{L}$. The appropriate value for this term should be chosen via cross-validation methods.